# Say My Name: a Model’s Bias Discovery Framework

### v1

@@ B.1 Ablations on the Bias Mining step
In this section we provide visualizations on the output distributions on the Waterbirds target when training a ResNet-50 on Waterbirds in the same setup as described in Sec. 4.1. Fig. 9 proposes visualizations of the output distributions for the bias-target aligned samples in blue (waterbirds and sea landscape), and bias-target conflicting samples in orange (waterbirds and ground landscape), in three different moments of the training: in the early stages (at $t=1$, on the left, where $t$ is the training epoch), at the chosen bias extraction time $t^{*}$ ($t=6$, at the center) and in the final stages ($t=10$, on the right). We recall that the entire bias extraction process happens directly on the training set $\mathcal{D}^{\text{train}}$ and does not require the employment of a validation/test set. We observe here that, at extraction time, there is an evident separation between bias-aligned and bias-conflicting samples, and we are able to confidently isolate the most biased among the conflicting samples (the orange distribution having a larger population below the random guess threshold). This does not hold in case the extraction time is delayed, with the bias-conflicting distribution not exhibiting a peak anymore.
B.2 Ablation on Bias Naming
B.2.1 Ablation on Text Embedders
@@
DistilRoberta [^6] https://huggingface.co/sentence-transformers/all-distilroberta-v1 Landbirds tree (0.26647)
Waterbirds marine (0.39470), ocean (0.36495), seagull (0.34526), coastal (0.34340), boat (0.33371), beach (0.33185), sea (0.31782), shoreline (0.30433), sandy (0.24345), waves (0.21502)
All-MiniLM-L12-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 Landbirds tree (0.34342), forest (0.31474), trees (0.30727), shrubs (0.28954), foliage (0.28264), vegetation (0.25404), branch (0.25149), leaves (0.21209)
Waterbirds sea (0.41955), ocean (0.37026), boat (0.35566), tide (0.34441), seagull (0.32581), flight (0.28337), waves (0.28000), coastal (0.27740), shoreline (0.25667), lake (0.23938), beach (0.23500), marine (0.22872)
MPNET [^8] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 Landbirds tree (0.38796), trees (0.32103), forest (0.30468), shrubs (0.25183), foliage (0.21930), branch (0.21547), bamboo (0.20619)
Waterbirds sea (0.39526), ocean (0.33487), beach (0.33385), shoreline (0.33176), waves (0.32562), lagoon (0.31923), coastal (0.30199), boat (0.26389), marine (0.24594), seagull (0.22533), water (0.20007)
NOMIC [^9] https://huggingface.co/nomic-ai/nomic-embed-text-v1 Landbirds small (0.24587), forest (0.21252), branch (0.20939)
Waterbirds sea (0.29738), ocean (0.28037), shoreline (0.27531), waves (0.27215), beach (0.25334), coastal (0.24178), tide (0.23727), seagull (0.22728), boat (0.22523), lagoon (0.21229), lake (0.20453), marine (0.20336)
AlBERT [^10] https://huggingface.co/sentence-transformers/paraphrase-albert-small-v2 Landbirds tree (0.39055), branch (0.34739), trees (0.34374), forest (0.32672), shrubs (0.29123), vegetation (0.28895), foliage (0.28274), bamboo (0.25158), left (0.21710), leaves (0.21549)
Waterbirds beach (0.41926), ocean (0.39030), sea (0.38636), shoreline (0.35808), waves (0.35194), tide (0.34131), coastal (0.33961), marine (0.30102), sandy (0.28278), midflight (0.27492), boat (0.27448), seagull (0.25469), water (0.24652), lagoon (0.22010), flight (0.20572)
@@
In this subsection, we are interested in tasting SaMyNa with more diverse models for the textual embedding, in an attempt to check the generality of the proposed approach. We provide, in Tab. 2, the results obtained with five other popular textual embedding models. From our results, we can clearly see that when employing any of the tested models we are able to find back the two typical biases from waterbirds. We observe though that bigger encoders provide higher similarity scores (MPNET and AlBERT), while smaller ones show lower scores, typically for the bias associated with landbirds. We can explain this to the higher complexity of capturing more diverse features associated with ground environments rather than maritime ones, requiring a higher capacity from the encoder. This pushes us to use, in an agnostic setup where we place ourselves in, to use a large generic encoder.
B.2.2 Ablation on $\bf f_{min}$
We propose here, in Tab. 3, the results we obtain for varying values for the hyperparameter $f_{\text{min}}$, i.e. the minimum frequency with which a keyword has to appear so that it can be considered as a possible output. What is possible to observe is that the higher $f_{\text{min}}$ the fewer and fewer keywords will be selected, and at some point, one class will be wrongfully marked as not holding a bias (with $f_{\text{min}}=0.60$). On the other hand, when not employing filtering at all ($f_{\text{min}}=0$), a lot of more fine-grained classes like coniferous or brownish arise. We find the chosen threshold ($f_{\text{min}}=0.15$) a fair compromise. As a sanity check, we also observe no change in the score for the remaining keywords.
@@ B.2.2 Ablation on $\bf f_{min}$
tree 0.36 sea 0.53
shrubs 0.35 ocean 0.51
forest 0.34 beach 0.45
branch 0.33 waters 0.43
deciduous 0.33 shoreline 0.42
foliage 0.33 shore 0.41
branches 0.32 sailing 0.40
twigs 0.32 aquatic 0.40
trees 0.32 harbor 0.39
stalks 0.31 tide 0.39
forested 0.31 ship 0.38
plants 0.30 pier 0.38
plant 0.30 coastal 0.37
wooded 0.30 beachfront 0.37
leafy 0.29 coastline 0.36
woodland 0.27 water 0.33
vegetation 0.26 sailboat 0.32
brownishgray 0.24 marine 0.32
coniferous 0.24 boat 0.31
greenery 0.23 submerged 0.31
leafless 0.23 pond 0.30
brownishgrey 0.23 lagoon 0.29
leaves 0.22 waves 0.28
grasses 0.22 vessel 0.28
brownish 0.21 swimming 0.28
stalk 0.21 motorboat 0.27
midflight 0.27
boathouse 0.27
lake 0.27
@@
wave 0.26
flying 0.26
boating 0.26
@@
river 0.25
reef 0.25
watermark 0.23
cranes 0.23
gliding 0.22
fish 0.22
seagulls 0.22
hull 0.20
flight 0.20
@@
tree 0.36 sea 0.53
shrubs 0.35 ocean 0.51
forest 0.34 beach 0.45
branch 0.33 shoreline 0.42
foliage 0.33 shore 0.41
branches 0.32 tide 0.39
trees 0.32 coastal 0.37
stalks 0.31 water 0.33
forested 0.31 marine 0.32
plants 0.30 boat 0.31
wooded 0.30 submerged 0.31
vegetation 0.26 lagoon 0.29
coniferous 0.24 waves 0.28
greenery 0.23 midflight 0.27
leaves 0.22 lake 0.27
river 0.25
flight 0.20
@@
tree 0.36 sea 0.53
shrubs 0.35 ocean 0.51
forest 0.34 beach 0.45
branch 0.33 shoreline 0.42
foliage 0.33 tide 0.39
trees 0.32 coastal 0.37
vegetation 0.26 water 0.33
greenery 0.23 marine 0.32
leaves 0.22 boat 0.31
lagoon 0.29
waves 0.28
midflight 0.27
lake 0.27
flight 0.20
@@
forest 0.34 ocean 0.51
branch 0.33 beach 0.45
trees 0.32 coastal 0.37
leaves 0.22 water 0.33
waves 0.28
midflight 0.27
lake 0.27
flight 0.20
@@
branch 0.33 water 0.33
waves 0.28
flight 0.20
@@
B.2.3 Ablation on $k$
We present here the ablation on $k$, that selects the cardinality of aligned and conflicting samples per class. Fig. 10 reports the study for 5 most occurring keywords in the cases under exam, while the full results are later reported in Table 12. In the general case, we observe that for lower values of $k$ the similarity score is in general lower, evidencing that the information extraction process is less accurate due to a general lack of information (and in general variety). This trend is particularly evident in more generic keywords like forest, trees, and beach. Finer-grained keywords like foliage and waves show a more irregular trend due to the specific sample selection. Overall we find that a fair compromise between performance and complexity is given by the intermediate $k=10$ for which some keywords like trees reach a high value standing constant for higher values of $k$. We highlight that maintaining $k$ at bay reduces the number of comparisons to perform, given that they grow quadratically.
B.2.4 Ablation on $\bf t_{min}$
@@
In this ablation study, we provide an example of the keywords resulting from SaMyNa when not employing any thresholding on the minimum similarity values for the found keywords. In Tab. 4, we provide the full output only for landbirds, due to the excessive length of the unfiltered output. We also provide the full list of the resulting keywords as a text file, available in the supplementary materials zip archive (keywords_ablation_tmin.txt). From an analysis of the emerging keywords, we observe three interesting intervals of values. High similarity values indicate those concepts that correlate well specifically with the learned class (filtered from the target class) and are those presented in the paper. Concepts whose similarity is close to zero are not correlated: indeed, we can find keywords like long, muted, and given that are neutral concepts. Interestingly, we can also identify concepts like background and environment that are super-classes of the two biases. This confirms that SaMyNa works properly since it puts itself in the best spot to best discriminate the two biases. Finally, the third region is for negatively correlated concepts (anti-correlated), where we easily find the concepts correlated with the other class (waterbirds) given that we are in a binary classification task.
Use of $t_{min}$ Keyword (value)
@@
$t_{min}=-\infty$ (Landbirds) tree (0.36349), shrubs (0.35333), forest (0.34249), branch (0.33068), foliage (0.32688), trees (0.31639), vegetation (0.25899), greenery (0.23251), leaves (0.22270), rounded (0.17975), bamboo (0.16404), green (0.16104), nature (0.14898), species (0.14680), brown (0.14534), small (0.13999), black (0.13964), gray (0.13468), natural (0.12627), habitat (0.12563), yellow (0.12454), bird (0.11301), patch (0.11042), perched (0.10776), soft (0.10717), neck (0.10514), lush (0.10440), slightly (0.09828), trunks (0.09826), standing (0.09586), color (0.09571), colors (0.09523), dark (0.09419), darker (0.09084), feet (0.09080), wildlife (0.08827), turned (0.08459), looking (0.07803), facing (0.07712), ecosystem (0.07514), palette (0.07354), lighter (0.07160), orange (0.07136), lighting (0.07126), position (0.06891), left (0.06699), markings (0.06692), positioned (0.06294), vibrant (0.05924), beak (0.05871), slender (0.05838), round (0.05730), side (0.05356), calm (0.05190), bright (0.04925), vertical (0.04921), pointed (0.04911), ground (0.04846), short (0.04838), red (0.04834), floor (0.04783), suggests (0.04622), blue (0.04544), tones (0.04530), passerine (0.04368), area (0.04230), plumage (0.04097), surrounding (0.04089), thin (0.03906), fallen (0.03622), gives (0.03523), birds (0.03506), main (0.03480), predominantly (0.03343), distinctive (0.03313), setting (0.03305), features (0.03300), giving (0.03169), found (0.03098), elements (0.03031), light (0.02905), type (0.02880), structures (0.02876), tall (0.02846), covered (0.02709), tail (0.02559), open (0.02540), feathers (0.02521), eyes (0.02402), appears (0.02250), suggesting (0.02136), indicating (0.02100), consists (0.02100), scattered (0.01934), wingtips (0.01928), gently (0.01406), subject (0.01349), might (0.01296), reflects (0.01232), possibly (0.01221), peaceful (0.01160), tranquil (0.01087), thriving (0.01059), sense (0.00703), presence (0.00641), tranquility (0.00393), tropical (0.00369), characterized (0.00365), suggest (0.00195), muted (0.00093), back (0.00062), seems (-0.00001), legs (-0.00070), near (-0.00084), towards (-0.00113), background (-0.00193), effect (-0.00199), eye (-0.00500), point (-0.00510), likely (-0.00543), providing (-0.00555), bare (-0.00591), given (-0.01045), environment (-0.01051), deep (-0.01359), visible (-0.01432), right (-0.01460), either (-0.01469), moment (-0.01475), could (-0.01496), long (-0.01563), dense (-0.01603), across (-0.01712), large (-0.01920), fully (-0.02127), focal (-0.02175), appear (-0.02263), creating (-0.02425), shows (-0.02666), healthy (-0.02720), cloudy (-0.02783), prominent (-0.02938), various (-0.03056), conditions (-0.03063), camera (-0.03170), photo (-0.03228), viewer (-0.03262), head (-0.03315), buildings (-0.03439), composition (-0.03512), overall (-0.03523), day (-0.03579), depicts (-0.03646), midst (-0.03986), daytime (-0.04005), focus (-0.04202), along (-0.04251), typical (-0.04274), taking (-0.04317), sunny (-0.04389), drawing (-0.04575), blurred (-0.04678), foreground (-0.04908), activity (-0.04970), contrast (-0.05024), wings (-0.05129), landscape (-0.05249), surroundings (-0.05353), packed (-0.05497), white (-0.05643), beauty (-0.05706), movement (-0.05968), weather (-0.06107), image (-0.06449), one (-0.06574), attention (-0.06792), captured (-0.06949), taken (-0.06966), extended (-0.06995), scene (-0.07093), view (-0.07824), clearly (-0.08422), atmosphere (-0.08640), mix (-0.08732), food (-0.08997), body (-0.09025), serene (-0.09172), similar (-0.09231), distance (-0.09281), backdrop (-0.09946), clouds (-0.10134), dynamic (-0.10504), surface (-0.10775), captures (-0.11025), overcast (-0.11541), world (-0.12448), sky (-0.12458), picturesque (-0.12616), horizon (-0.14756), clear (-0.15543), sandy (-0.15546), landing (-0.17462), seagull (-0.19202), flight (-0.20084), lake (-0.26879), midflight (-0.27087), waves (-0.28129), lagoon (-0.28892), boat (-0.31066), marine (-0.31538), water (-0.32785), coastal (-0.37410), tide (-0.38738), shoreline (-0.41605), beach (-0.45110), ocean (-0.51175), sea (-0.52685)

### v2

@@ C.1 Ablations on the Bias Mining step
In this section, we provide visualizations on the output distributions on the Waterbirds target when training a ResNet-50 on Waterbirds in the same setup as described in Sec. 4.1. Fig. 11 proposes visualizations of the output distributions for the bias-target aligned samples in blue (waterbirds and sea landscape), and bias-target conflicting samples in orange (waterbirds and ground landscape), in three different moments of the training: in the early stages (at $t=1$, on the left, where $t$ is the training epoch), at the chosen bias extraction time $t^{*}$ ($t=6$, at the center) and in the final stages ($t=10$, on the right). We recall that the entire bias extraction process happens directly on the training set $\mathcal{D}^{\text{train}}$ and does not require the employment of a validation/test set. We observe here that, at extraction time, there is an evident separation between bias-aligned and bias-conflicting samples, and we are able to confidently isolate the most biased among the conflicting samples (the orange distribution having a larger population below the random guess threshold). This does not hold in case the extraction time is delayed, with the bias-conflicting distribution not exhibiting a peak anymore.
C.2 Ablation on Bias Naming
C.2.1 Ablation of Eq. 4
To demonstrate the effectiveness of Eq. 4 in filtering out the shared information across all dataset we did an ablation study on the CelebA dataset. Fig. 12 shows the top-9 keywords for each class on CelebA when Eq. 4 is removed from our method. As can be seen, the ranking is full of irrelevant keywords that can be used to describe just about any image in the dataset, like portrait, photo, image, and headshot.
Figure 12: Similarity scores for the CelebA dataset when ablating Eq. 4 from our method.
C.2.2 Ablation on Text Embedders
@@
DistilRoberta Landbirds tree (0.22477)
Waterbirds ocean (0.41353), shore (0.37904), beach (0.37184), boat (0.35631), water (0.28810), river (0.28312), waves (0.28209), sandy (0.25046), seagull (0.23950), lake (0.22972), across (0.21785), foam (0.20274)
All-MiniLM-L12-v2 Landbirds tree (0.35688), forest (0.33959), trees (0.31249), foliage (0.27917), branches (0.25548), vegetation (0.25477), branch (0.25477), leaves (0.21215)
Waterbirds ocean (0.40874), boat (0.35452), shore (0.31332), waves (0.27757), beach (0.27478), seagull (0.23625), lake (0.22088)
MPNET Landbirds tree (0.38503), forest (0.31970), trees (0.31858), foliage (0.26440), branches (0.25911), branch (0.23637), bird (0.22693), bamboo (0.22661), perched (0.20324)
Waterbirds waves (0.34030), ocean (0.33423), beach (0.31522), shore (0.26980), water (0.22304)
NOMIC Landbirds forest (0.22448), branches (0.21085)
Waterbirds waves (0.30429), ocean (0.29857), beach (0.26858), water (0.23373), boat (0.21911), shore (0.21033), lake (0.20263)
AlBERT Landbirds trees (0.32898), tree (0.32710), foliage (0.31355), forest (0.30978), branch (0.29175), branches (0.29141), vegetation (0.28953), bamboo (0.27675), stalks (0.23657)
Waterbirds ocean (0.49492), beach (0.47047), waves (0.42517), water (0.38388), shore (0.35404), boat (0.29732), foam (0.26902), lake (0.26790), sandy (0.25054), surface (0.24005), midflight (0.21686), river (0.21444)
@@
In this subsection, we are interested in testing SaMyNa with more diverse models for the textual embedding, in an attempt to check the generality of the proposed approach. We provide, in Tab. 9, the results obtained with five other popular textual embedding models. From our results, we can clearly see that when employing any of the tested models we are able to find back the two typical biases from waterbirds. All models produce similar outputs, except for DistilRoberta and NOMIC that produce very few, albeit correct, keywords for Landbirds.
C.2.3 Ablation on $\bf f_{min}$
We propose here, in Table˜10, the results we obtain by varying the value of the hyperparameter $f_{\text{min}}$, i.e. the minimum fraction of captions from examples of the same predicted class in which a keyword must appear so that it can be considered as a possible output. As $f_{\text{min}}$ increases, each new set of selected keywords is a subset of the previous one, with all keywords maintaining identical scores. When $f_{\text{min}}$ is too high, one class will be wrongfully marked as not holding a bias (with $f_{\text{min}}=0.60$). On the other hand, when not employing filtering at all ($f_{\text{min}}=0$), a lot of more fine-grained classes like coniferous or wilderness arise. We chose a low threshold ($f_{\text{min}}=0.15$) to filter out just the least frequent words that may come from mistakes of the captioner or from details that appear in just a few images. Filtering out too many words with this method may remove important synonyms from the ranking, so we avoid it.
@@ C.2.3 Ablation on $\bf f_{min}$
shrubs 0.35 sea 0.54
forest 0.33 ocean 0.54
foliage 0.33 aquatic 0.44
twigs 0.33 beach 0.43
tree 0.33 harbor 0.40
deciduous 0.32 ships 0.39
stalks 0.31 shore 0.39
branch 0.31 water 0.39
forested 0.31 shoreline 0.38
trees 0.30 pier 0.37
branches 0.30 ship 0.37
bushes 0.30 coastal 0.35
wooded 0.29 waves 0.33
vegetation 0.27 sailboat 0.32
weeds 0.26 coastline 0.32
grasses 0.24 sail 0.30
greenery 0.23 wave 0.30
wilderness 0.22 vessel 0.30
stalk 0.22 floating 0.29
coniferous 0.22 boat 0.29
garden 0.21 speedboat 0.29
bark 0.20 wetsuit 0.29
grayishbrown 0.20 pond 0.29
swimming 0.28
lagoon 0.28
bay 0.27
splash 0.27
river 0.27
surfboard 0.27
@@
surfer 0.26
glides 0.25
surfing 0.25
@@
pool 0.25
midflight 0.24
lake 0.24
cranes 0.22
flying 0.22
sand 0.21
watermark 0.20
@@
shrubs 0.35 ocean 0.54
forest 0.33 beach 0.43
foliage 0.33 shore 0.39
tree 0.33 water 0.39
deciduous 0.32 shoreline 0.38
stalks 0.31 coastal 0.35
branch 0.31 waves 0.33
trees 0.30 wave 0.30
branches 0.30 boat 0.29
vegetation 0.27 river 0.27
grasses 0.24 midflight 0.24
lake 0.24
flying 0.22
watermark 0.20
@@
forest 0.33 ocean 0.54
foliage 0.33 beach 0.43
tree 0.33 shore 0.39
stalks 0.31 water 0.39
branch 0.31 waves 0.33
trees 0.30 boat 0.29
branches 0.30 river 0.27
vegetation 0.27 midflight 0.24
lake 0.24
flying 0.22
@@
forest 0.33 ocean 0.54
branch 0.31 water 0.39
trees 0.30 waves 0.33
@@
trees 0.30 water 0.39
@@
C.2.4 Ablation on $\bf t_{sim}$
@@
In this ablation study, we provide an example of the keywords resulting from SaMyNa when not employing any thresholding on the minimum similarity values for the found keywords. In Table˜11, we provide the full output only for landbirds, due to the excessive length of the unfiltered output. However, this is not an issue because SaMyNa, when applied to binary classification, possesses a mathematical property that guarantees symmetrical keyword rankings for the two classes. Consequently, the keyword ranking for waterbirds is simply the inverse of the ranking for landbirds (the cosine similarity scores are the same but with opposite sign). Nevertheless, we provide the full list of the resulting keywords as a text file, available in the supplementary materials zip archive (keywords_ablation_tsim.txt). From an analysis of the emerging keywords, we observe three interesting intervals of values. High similarity values indicate those concepts that correlate well specifically with the learned class and are those presented in the paper. Concepts whose similarity is close to zero are not correlated: indeed, we can find keywords like possibly, various, and open that are neutral concepts. Interestingly, we can also identify concepts like background and environment that are super-classes of the two biases. This confirms that SaMyNa works properly since it puts itself in the best spot to best discriminate the two biases. Finally, the third region is for negatively correlated concepts (anti-correlated), where we easily find the concepts correlated with the other class (waterbirds) given that we are in a binary classification task.
Use of $t_{\text{sim}}$ Keyword (value)
@@
$t_{\text{sim}}=-1$ (Landbirds) forest (0.33454), foliage (0.33412), tree (0.32720), stalks (0.31389), branch (0.31242), trees (0.29919), branches (0.29649), vegetation (0.27349), leaves (0.19398), bamboo (0.16775), brown (0.16085), small (0.15988), black (0.15947), stands (0.15171), green (0.15048), gray (0.14847), nature (0.14642), yellow (0.14465), perched (0.14381), habitat (0.14101), positioned (0.12865), corner (0.12602), position (0.12585), style (0.12585), standing (0.12558), color (0.12359), bird (0.12310), natural (0.12179), webbed (0.11827), naturalistic (0.11359), lush (0.11342), darker (0.11082), dark (0.10989), colors (0.10947), features (0.10804), slightly (0.10770), texts (0.10516), photography (0.10478), markings (0.10345), lighting (0.10088), plumage (0.09747), shades (0.09703), setting (0.09479), camera (0.08995), slender (0.08957), soft (0.08622), feathers (0.08274), beak (0.08256), pointed (0.08064), facing (0.08016), center (0.07854), feet (0.07604), area (0.06895), vibrant (0.06787), short (0.06720), frame (0.06472), depicts (0.06152), appears (0.06135), central (0.06021), background (0.06013), ground (0.05912), creating (0.05908), photograph (0.05901), elements (0.05896), side (0.05860), located (0.05815), main (0.05688), subject (0.05661), top (0.05637), snapshot (0.05501), looking (0.05431), marks (0.05423), focal (0.05365), eyes (0.05237), suggests (0.05027), composite (0.04975), eye (0.04772), lighter (0.04710), presence (0.04675), create (0.04601), tall (0.04574), turned (0.04134), perspective (0.03960), indicating (0.03954), calm (0.03905), predominantly (0.03879), suggesting (0.03870), path (0.03797), thin (0.03732), light (0.03667), moment (0.03628), left (0.03577), touch (0.03549), surrounding (0.03547), wide (0.03536), lower (0.03233), featuring (0.03144), distinctive (0.03032), landscape (0.02956), onto (0.02917), red (0.02857), consists (0.02824), foreground (0.02821), bottom (0.02585), scale (0.02508), sense (0.02173), context (0.02116), focus (0.02088), possibly (0.02039), capturing (0.01853), around (0.01757), legs (0.01733), visible (0.01702), day (0.01428), visually (0.01370), composition (0.01359), point (0.01284), scene (0.01166), might (0.01119), making (0.01009), blue (0.01004), distinguishing (0.01002), reflecting (0.00721), tail (0.00713), amidst (0.00643), open (0.00581), objects (0.00573), providing (0.00562), quality (0.00505), towards (0.00415), image (0.00329), taking (0.00250), captured (0.00037), buildings (-0.00146), viewer (-0.00204), peaceful (-0.00231), taken (-0.00423), wings (-0.00644), various (-0.00649), environment (-0.00728), movement (-0.00924), view (-0.00953), head (-0.01000), scattered (-0.01045), provide (-0.01047), overall (-0.01051), fully (-0.01158), detail (-0.01264), captures (-0.01321), near (-0.01394), likely (-0.01408), tranquil (-0.01455), appealing (-0.01471), drawing (-0.01530), body (-0.01555), expanse (-0.01751), could (-0.01791), backdrop (-0.01856), large (-0.02118), immediately (-0.02224), white (-0.02302), right (-0.02431), adding (-0.02552), realistic (-0.03162), deep (-0.03255), surroundings (-0.03269), otherwise (-0.03385), attention (-0.03709), beautifully (-0.03766), relative (-0.03802), beautiful (-0.03860), sunny (-0.04031), tranquility (-0.04087), freedom (-0.04306), contrast (-0.04520), also (-0.04593), sky (-0.04655), across (-0.04660), blend (-0.04802), extended (-0.04905), depth (-0.05085), juxtaposes (-0.05146), dense (-0.05250), adds (-0.06059), two (-0.07112), clouds (-0.07306), dynamic (-0.07317), spread (-0.07693), contrasting (-0.07756), surface (-0.07825), harmonious (-0.08196), serene (-0.08365), world (-0.10627), distance (-0.10900), moving (-0.11027), mix (-0.11326), filled (-0.12270), soars (-0.13097), overcast (-0.14479), flight (-0.14884), sandy (-0.14945), clear (-0.16600), seagull (-0.18060), foam (-0.19652), flying (-0.21832), lake (-0.24355), midflight (-0.24404), river (-0.26848), boat (-0.29352), waves (-0.33462), water (-0.38571), shore (-0.38639), beach (-0.43417), ocean (-0.53961)